OpenStudy (anonymous):

Show an equation and a solution. How much pure antifreeze must be added to 12 L of a 40% solution to obtain a 60% solution. (Remember that pure antifreeze is 100% = 1.) Question 7 answers

OpenStudy (anonymous):

help @Meepi

OpenStudy (anonymous):

@ryan123345

OpenStudy (anonymous):

help

OpenStudy (anonymous):

Write an equation and a solution for the problem. A health food store sells oatmeal for $3.60 per pound and bran flakes for $4.80 per pound. How many pounds of each should be used to get a mixture of 30 pounds that sells for $4.00 a pound? and that

OpenStudy (anonymous):

You have a solution that currently contains 0.4 * 12 = 4.8 liters of antifreeze (40%), and you want to get a 60% solution by adding 100% antifreeze if you add x liters of antifreeze, you will have a total of 4.8 + x liters of antifreeze, in 12 + x liters. You want this to be a 60% antifreeze per liter solution so: \[\frac{4.8 + x}{12 + x} = \frac{6}{10}\] This is the last one I can explain I really have to go now :)